### v1

@@ 7 Conclusion and future outlook
$\frac{\partial e^{n}}{\partial t}+\frac{\partial}{\partial x}\big{(}\alpha^{n}(x)e^{n}\big{)}$ $=0,\quad(x,t)\in(-1,1)\times(t_{n},t_{n+1}],$ (SM29)
where $\quad e^{n}(x,t_{n})=e^{n-1}(x,t_{n})$. The PDE (SM29) arises from considering (cons-lin) over the local time interval $t\in(t_{n},t_{n+1}]$ and freezing the time-dependence of the wave-speed on this interval at the point $t_{n}+\delta t\vartheta$ [^9] Our analysis is agnostic to the choice of $\vartheta\in[0,1]$; however, when the results of this analysis are used heuristically to create an MGRIT coarse-grid operator, we find the choice of $\vartheta$ is extremely important in some cases. For example, in the MGRIT solution of the linearized systems arising from a 3rd-order WENO discretization of the Buckley–Leverett test problems from the main paper. We now outline the discretization for this problem, which tracks closely with the discretization outlined in Section 2; however, they are some key differences.
Integrating (SM29) over the discretized spatial domain gives the exact relation for $t\in(t_{n},t_{n+1}]$
$\frac{\operatorname{d}\!\bar{e}^{n}_{i}(t)}{\operatorname{d}\!t}=-\frac{f^{n}(x_{i+1/2},e(x_{i+1/2}))-f^{n}(x_{i-1/2},e(x_{i-1/2}))}{h},\quad i\in\{1,\ldots,n_{x}\},$ (SM30)
@@
Next, the physical fluxes in (SM30) are approximated with numerical fluxes:
@@
$\frac{\operatorname{d}\!\bar{e}_{i}^{n}}{\operatorname{d}\!t}\approx-\frac{\widehat{f}^{n}_{i+1/2}-\widehat{f}^{n}_{i-1/2}}{h}=:\big{(}\widehat{L}^{n}\bar{\bm{e}}^{n}\big{)}_{i},\quad i\in\{1,\ldots,n_{x}\},$ (SM31)
where $\bar{\bm{e}}^{n}=\big{(}\bar{e}_{1}^{n},\ldots,\bar{e}_{n_{x}}^{n}\big{)}^{\top}\in\mathbb{R}^{n_{x}}$. The numerical flux $\widehat{f}^{n}_{i+1/2}$ introduced here is of the LF type for a linear problem given by (12). For the time being, we do not want to make any assumptions about the numerical flux, including how $e^{n}(x_{i+1/2})$ is reconstructed. [^10] Later we consider several different reconstruction techniques including standard polynomial reconstruction and Picard-linearized WENO reconstructions. Therefore our goal is to develop a general error estimate that is agnostic to specifics of the reconstruction. That being said, the analysis requires us to make the following assumption. {assumption}[Error for general linear spatial discretization] Assume that the numerical flux function $\widehat{f}^{n}_{i+1/2}=\widehat{f}^{n}_{i+1/2}(\bar{\bm{e}}^{n})\approx\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})$ introduced in (SM31) is linear in $\bar{\bm{e}}^{n}$. Further, assume that the numerical flux satisfies
$\widehat{f}^{n}_{i+1/2}(\bar{\bm{e}}^{n})=\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})+\widehat{E}^{n}_{i+1/2}(\bar{\bm{e}}^{n}),$ (SM32)
@@
Defining the linear difference operator ^$\widehat{{\cal E}}^{n}\colon\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ such that (^$\big{(}\widehat{{\cal E}}^{n}\bar{\bm{e}}^{n}\big{)}_{i}=\frac{\widehat{E}^{n}_{i+1/2}-\widehat{E}^{n}_{i-1/2}}{h}$, from the above relation, it follows that the right-hand side of (SM30) can be expressed as a linear function of $\bar{\bm{e}}^{n}$ as follows:
$-\frac{\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})-\alpha^{n}(x_{i-1/2})e^{n}(x_{i-1/2})}{h}=\big{[}\big{(}\widehat{L}^{n}+\widehat{{\cal E}}^{n}\big{)}\bar{\bm{e}}^{n}\big{]}_{i},$ (SM33)
with $\widehat{L}^{n}$ the spatial discretization in (SM31). Once specific details of the reconstruction are specified, estimates can be made for the error term $\widehat{E}_{i+1/2}$ (see, e.g., Lemma SM5.3). Since the flux is assumed linear in $\bar{\bm{e}}^{n}$, it follows that the spatial discretization operator $\widehat{L}^{n}$ from (SM31) is also assumed to be linear.
Next, the ODEs (SM31) are approximately advanced forwards in time with an ERK method resulting in
$\bar{\bm{e}}^{n+1}=\Phi^{n}\bar{\bm{e}}^{n},$ (SM34)
with $\Phi^{n}$ the time-stepping operator, and $\bar{\bm{e}}^{n}\approx\bar{\bm{e}}(t_{n})$. Since the ODEs (SM31) are linear and autonomous, the time-stepping operator can be written as
$\Phi^{n}=R\big{(}\delta t\widehat{L}^{n}\big{)}=\exp\big{(}\delta t\widehat{L}^{n}\big{)}+\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+1}+{\cal O}\big{(}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+2}\big{)},$ (SM35)
@@
SM4.2 Truncation error estimate
We now develop a truncation error estimate for the previously described discretization of the integral conservation law (SM30). First Lemma SM4.1 presents an exact solution to (SM30), and then Lemma SM4.3 gives a truncation error estimate for the discretization of it.
Lemma SM4.1 (Exact solution of integral form of conservation law).
Assume that Footnote 10 holds, and let $\bar{\bm{e}}^{n}(t)$ be the exact solution of the local conservation law (SM30) at time $t\in[t_{n},t_{n+1}]$. Then,
$\bar{\bm{e}}^{n}(t)=\exp\big{[}(t-t_{n})\big{(}\widehat{L}^{n}+{\cal E}^{n}\big{)}\big{]}\bar{\bm{e}}^{n}(t_{n}),\quad t\in[t_{n},t_{n}+\delta t],$ (SM36)
@@
To see that (SM36) indeed solves (SM30), consider the $i$th component of (SM36), differentiating both sides with respect to $t$ and then invoking the expression for $\widehat{L}^{n}+\widehat{{\cal E}}^{n}$ from Footnote 10:
$\frac{\operatorname{d}\!\bar{e}_{i}^{n}}{\operatorname{d}\!t}$ $=\Big{[}\big{(}\widehat{L}^{n}+\widehat{{\cal E}}^{n}\big{)}\exp\big{[}(t-t_{n})\big{(}\widehat{L}^{n}+{\cal E}^{n}\big{)}\big{]}\bar{\bm{e}}^{n}(t_{n})\Big{]}_{i},$ (SM37)
$=\big{[}\big{(}\widehat{L}^{n}+\widehat{{\cal E}}^{n}\big{)}\bar{\bm{e}}^{n}\big{]}_{i},$ (SM38)
$=-\frac{\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})-\alpha^{n}(x_{i-1/2})e^{n}(x_{i-1/2})}{h}.$ (SM39)
@@
We are now ready to present an error estimate for the fully discrete scheme (SM34). The steps in this proof track closely with those used in the proof of [16, Lemma 3], in which we developed a truncation error estimates for non-conservative, FD, Runge-Kutta discretizations of constant wave-speed linear advection problems.
Lemma SM4.3 (Truncation error estimate for $\Phi^{n}$).
@@
Assume that Footnote 10 holds, and let $\bar{\bm{e}}^{n}(t)$ be the exact solution of the local conservation law (SM30) at time $t\in[t_{n},t_{n+1}]$. Assume this function is sufficiently smooth. Then, the local truncation error of the one-step discretization (SM34) at time $t_{n+1}$ can be written as
$\bm{\tau}^{n}$ $:=\bar{\bm{e}}^{n}(t_{n+1})-\Phi^{n}\bar{\bm{e}}^{n}(t_{n}),$ (SM40)
$=\Big{(}-\delta t\widehat{{\cal E}}^{n}+\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+1}\Big{)}\bar{\bm{e}}^{n}(t_{n+1})+{\cal O}\Big{(}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+2},\big{(}\delta t\widehat{{\cal E}}^{n}\big{)}^{2}\Big{)},$ (SM41)
@@
in which $\widehat{L}^{n}$ is the spatial discretization introduced in (SM31), and ^$\widehat{{\cal E}}^{n}$, as defined in Footnote 10, is the associated spatial discretization error.
Proof SM4.4.
Using (SM36), the solution to (SM30) at time $t=t_{n}$ may be written as
$\bar{\bm{e}}^{n}(t_{n})=\exp\big{(}-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big{)}\bar{\bm{e}}^{n}(t_{n+1}).$ (SM42)
Plugging this into definition (SM40) gives
$\bm{\tau}^{n}=\Big{[}I-\Phi^{n}\exp\big{(}-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big{)}\Big{]}\bar{\bm{e}}^{n}(t_{n+1}).$ (SM43)
@@
Next, plug in the series expansion for $\Phi^{n}$ given in (SM35) and simplify to give (abusing notation by dropping the vector $\bar{\bm{e}}^{n}(t_{n+1})$)
$\begin{split}&I-\Phi^{n}\exp\big{(}-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big{)}\\ &=I-\Big{[}\exp\big{(}\delta t\widehat{L}^{n}\big{)}+\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+1}+{\cal O}\Big{(}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+2}\Big{)}\Big{]}\exp\big{(}-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big{)},\end{split}$ (SM44)
$=\Big{(}\delta t\widehat{{\cal E}}^{n}-\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+1}\Big{)}+{\cal O}\Big{(}\big{(}\delta t\widehat{L}^{n}\big{)}^{q+2},\big{(}\delta t\widehat{{\cal E}}^{n}\big{)}^{2}\Big{)}.$ (SM45)
Plugging this final expressing into (SM43) concludes the proof.
@@
From Lemma SM4.3 we want to develop a truncation error estimate for the ideal coarse-grid operator $\Psi_{\rm ideal}^{n,m\delta t}:=\prod_{j=0}^{m-1}\Phi^{n+j}$, i.e., $m$ fine-grid steps across the coarse-grid interval $t\in[t_{n},t_{n+m}]$. This can be developed rigorously using Taylor-series-type argument, but this is not necessary for our purposes since this analysis is heuristic anyway. Considering the form of (SM41) we suppose that the ideal coarse-grid operator satisfies the approximation
@@
$\begin{split}&\bar{\bm{e}}^{n+m-1}(t_{n+m})-\Psi_{\rm ideal}^{n,m\delta t}\bar{\bm{e}}^{n}(t_{n})\approx\\ &\hskip 17.22217pt\Bigg{[}\sum_{j=0}^{m-1}\Big{(}\delta t\widehat{{\cal E}}^{n+j}-\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n+j}\big{)}^{q+1}\Big{)}\Bigg{]}\bar{\bm{e}}^{n+m-1}(t_{n+m}).\end{split}$ (SM46)
SM4.3 Error estimates for reconstructions
@@
We conclude Section SM4 with the following lemma: An error estimate for the left-shifted polynomial reconstructions that are described Section 2.2 and Section SM1.
@@
Lemma SM4.5 (Error estimate for shifted linear reconstruction using $k$ cells).
@@
$\big{(}\widetilde{R}^{\ell}\bar{\bm{u}}\big{)}_{i}$ $=u(x_{i-1/2})-\frac{(-h)^{k}\widetilde{\zeta}^{\ell}}{(k+1)!}\frac{\operatorname{d}\!^{k}u}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{i,\ell}(x_{i-1/2})},\quad\widetilde{\zeta}^{\ell}=(-1)^{\ell}\ell!(k-\ell)!,$ (SM47)
$\big{(}R^{\ell}\bar{\bm{u}}\big{)}_{i}$ $=u(x_{i+1/2})-\frac{(-h)^{k}\zeta^{\ell}}{(k+1)!}\frac{\operatorname{d}\!^{k}u}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{i,\ell}(x_{i+1/2})},\quad\zeta^{\ell}=(-1)^{\ell+1}(\ell+1)!(k-\ell-1)!,$ (SM48)
@@
Proof SM4.6.
Because the reconstruction polynomial is the derivative of an interpolation polynomial (see Section SM1), this proof more-or-less follows the same steps as one used to derive the error for finite-difference approximations to a first derivative (these too are derivatives of interpolation polynomials); see e.g., [35, Section 8.1] or [16, Lemma 2].
Recall that for fixed $\ell$ that $\big{(}\widetilde{R}^{\ell}\bar{\bm{u}}\big{)}_{i}=q_{i}^{\ell}(x_{i-1/2})$ and $\big{(}R^{\ell}\bar{\bm{u}}\big{)}_{i}=q_{i}^{\ell}(x_{i+1/2})$, where $q_{i}^{\ell}(x)=\operatorname{d}\!Q_{i}^{\ell}/\operatorname{d}\!x$ is a degree at most $k-1$ polynomial recovering the cell averages of $e(x)$ over $k\geq 1$ cells $\{{\cal I}_{i-\ell},\ldots,{\cal I}_{i-\ell+k-1}\}$. Moreover, $Q_{i}^{\ell}$ is the unique polynomial of degree at most $k$ that interpolates $U(x):=\int^{x}u(\xi)\operatorname{d}\!\xi$ at the $k+1$ interfaces of the cells in the above stencil: $Q_{i}^{\ell}(x_{j+1/2})=U(x_{j+1/2})$ at points $\{x_{j+1/2}\}_{j=i-\ell-1}^{i-\ell-1-k}$.
@@
$U(x)-Q^{\ell}(x)=\frac{K(x)}{(k+1)!}\frac{\operatorname{d}\!^{k+1}U}{\operatorname{d}\!x^{k+1}}\bigg{|}_{\xi_{\ell}(x)},\quad K(x):=\prod\limits_{j=-\ell-1}^{-\ell-1+k}(x-x_{j+1/2}),$ (SM49)
with $\xi_{\ell}(x)$ some (unknown) point in the interior of the interpolation stencil, $\xi_{\ell}(x)\in(x_{i-\ell-1/2},x_{i-\ell+k-1/2})$. Since $\operatorname{d}\!U/\operatorname{d}\!x=u$ and we desire an estimate for $u$, we differentiate (SM49) to get
$u(x)-q^{\ell}(x)=\frac{1}{(k+1)!}\frac{\operatorname{d}\!K}{\operatorname{d}\!x}\bigg{|}_{x}\frac{\operatorname{d}\!^{k}u}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{\ell}(x)}+\frac{K(x)}{(k+1)!}\frac{\operatorname{d}\!^{k+1}u}{\operatorname{d}\!x^{k+1}}\bigg{|}_{\xi_{\ell}(x)}\frac{\operatorname{d}\!\xi_{\ell}}{\operatorname{d}\!x}\bigg{|}_{x}.$ (SM50)
To arrive at the estimates (SM47) and (SM48), all that is required is to evaluate (SM50) at $x=x_{-1/2}$ and $x=x_{1/2}$, respectively. To this end, notice from (SM49) that $K(x_{\pm 1/2})=0$, and, thus, the second added on the right-hand side of (SM50) vanishes when the estimate is evaluated at $x_{\pm 1/2}$. Next, from (SM49) we have that
$\frac{\operatorname{d}\!K}{\operatorname{d}\!x}=\sum\limits_{s=-\ell-1}^{-\ell-1+k}\hskip 8.61108pt\prod\limits_{\begin{subarray}{c}j=-\ell-1\\ \neq s\end{subarray}}^{-\ell-1+k}(x-x_{j+1/2}).$ (SM53)
It is not difficult to show that $\operatorname{d}\!K(x_{-1/2})/\operatorname{d}\!x=(-h)^{k}\widetilde{\zeta}^{\ell}$, and $\operatorname{d}\!K(x_{1/2})/\operatorname{d}\!x=(-h)^{k}{\zeta}_{\ell}$. Evaluating (SM50) at $x_{\pm 1/2}$, and plugging in the above values concludes the proof.
SM5 MGRIT solution of linear conservation laws using standard, linear method-of-lines discretization

### v2

@@ Acknowledgments
(SM41) $\frac{\partial e^{n}}{\partial t}+\frac{\partial}{\partial x}\big(\alpha^{n}(x)e^{n}\big)$ $=0,\quad(x,t)\in(-1,1)\times(t_{n},t_{n+1}],$
where $\quad e^{n}(x,t_{n})=e^{n-1}(x,t_{n})$. The PDE (SM41) arises from considering (cons-lin) over the local time interval $t\in(t_{n},t_{n+1}]$ and freezing the time-dependence of the wave-speed on this interval at the point $t_{n}+\delta t\vartheta$. [^9] Our analysis is agnostic to the choice of $\vartheta\in[0,1]$; however, when the results of this analysis are used heuristically to create an MGRIT coarse-grid operator, we find the choice of $\vartheta$ is extremely important in some cases. For example, in the MGRIT solution of the linearized systems arising from a 3rd-order WENO discretization of the Buckley–Leverett test problems from the main paper. We now outline the discretization for this problem, which tracks closely with the discretization outlined in Section 2; however, they are some key differences.
Integrating (SM41) over the discretized spatial domain gives the exact relation for $t\in(t_{n},t_{n+1}]$
(SM42) $\frac{\operatorname{d}\!\bar{e}^{n}_{i}(t)}{\operatorname{d}\!t}=-\frac{f^{n}(x_{i+1/2},e(x_{i+1/2}))-f^{n}(x_{i-1/2},e(x_{i-1/2}))}{h},\quad i\in\{1,\ldots,n_{x}\},$
@@
Next, the physical fluxes in (SM42) are approximated with numerical fluxes:
@@
(SM43) $\frac{\operatorname{d}\!\bar{e}_{i}^{n}}{\operatorname{d}\!t}\approx-\frac{\widehat{f}^{n}_{i+1/2}-\widehat{f}^{n}_{i-1/2}}{h}=:\big(\widehat{L}^{n}\bar{\bm{e}}^{n}\big)_{i},\quad i\in\{1,\ldots,n_{x}\},$
where $\bar{\bm{e}}^{n}=\big(\bar{e}_{1}^{n},\ldots,\bar{e}_{n_{x}}^{n}\big)^{\top}\in\mathbb{R}^{n_{x}}$. The numerical flux $\widehat{f}^{n}_{i+1/2}$ introduced here is of the LF type for a linear problem given by (10). For the time being, we do not want to make any assumptions about the numerical flux, including how $e^{n}(x_{i+1/2})$ is reconstructed. [^10] Later we consider several different reconstruction techniques including standard polynomial reconstruction and Picard-linearized WENO reconstructions. Therefore our goal is to develop a general error estimate that is agnostic to specifics of the reconstruction. That being said, the analysis requires us to make the following assumption. {assumption}[Error for general linear spatial discretization] Assume that the numerical flux function $\widehat{f}^{n}_{i+1/2}=\widehat{f}^{n}_{i+1/2}(\bar{\bm{e}}^{n})\approx\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})$ introduced in (SM43) is linear in $\bar{\bm{e}}^{n}$. Further, assume that the numerical flux satisfies
(SM44) $\widehat{f}^{n}_{i+1/2}(\bar{\bm{e}}^{n})=\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})+\widehat{E}^{n}_{i+1/2}(\bar{\bm{e}}^{n}),$
@@
Defining the linear difference operator $\widehat{{\cal E}}^{n}\colon\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ such that $\big(\widehat{{\cal E}}^{n}\bar{\bm{e}}^{n}\big)_{i}=\frac{\widehat{E}^{n}_{i+1/2}-\widehat{E}^{n}_{i-1/2}}{h}$, from the above relation, it follows that the right-hand side of (SM42) can be expressed as a linear function of $\bar{\bm{e}}^{n}$ as follows:
(SM45) $-\frac{\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})-\alpha^{n}(x_{i-1/2})e^{n}(x_{i-1/2})}{h}=\big[\big(\widehat{L}^{n}+\widehat{{\cal E}}^{n}\big)\bar{\bm{e}}^{n}\big]_{i},$
with $\widehat{L}^{n}$ the spatial discretization in (SM43). Once specific details of the reconstruction are specified, estimates can be made for the error term $\widehat{E}_{i+1/2}$ (see, e.g., Lemma SM7.3). Since the flux is assumed linear in $\bar{\bm{e}}^{n}$, it follows that the spatial discretization operator $\widehat{L}^{n}$ from (SM43) is also assumed to be linear.
Next, the ODEs (SM43) are approximately advanced forwards in time with an ERK method resulting in
(SM46) $\bar{\bm{e}}^{n+1}=\Phi^{n}\bar{\bm{e}}^{n},$
with $\Phi^{n}$ the time-stepping operator, and $\bar{\bm{e}}^{n}\approx\bar{\bm{e}}(t_{n})$. Since the ODEs (SM43) are linear and autonomous, the time-stepping operator can be written as
(SM47) $\Phi^{n}=R\big(\delta t\widehat{L}^{n}\big)=\exp\big(\delta t\widehat{L}^{n}\big)+\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n}\big)^{q+1}+{\cal O}\big(\big(\delta t\widehat{L}^{n}\big)^{q+2}\big),$
@@
SM6.2 Truncation error estimate
We now develop a truncation error estimate for the previously described discretization of the integral conservation law (SM42). First Lemma SM6.1 presents an exact solution to (SM42), and then Lemma SM6.3 gives a truncation error estimate for the discretization of it.
Lemma SM6.1 (Exact solution of integral form of conservation law).
Assume that Footnote 10 holds, and let $\bar{\bm{e}}^{n}(t)$ be the exact solution of the local conservation law (SM42) at time $t\in[t_{n},t_{n+1}]$. Then,
(SM48) $\bar{\bm{e}}^{n}(t)=\exp\big[(t-t_{n})\big(\widehat{L}^{n}+{\cal E}^{n}\big)\big]\bar{\bm{e}}^{n}(t_{n}),\quad t\in[t_{n},t_{n}+\delta t],$
@@
To see that (SM48) indeed solves (SM42), consider the $i$th component of (SM48), differentiating both sides with respect to $t$ and then invoking the expression for $\widehat{L}^{n}+\widehat{{\cal E}}^{n}$ from Footnote 10:
(SM49) $\frac{\operatorname{d}\!\bar{e}_{i}^{n}}{\operatorname{d}\!t}$ $=\Big[\big(\widehat{L}^{n}+\widehat{{\cal E}}^{n}\big)\exp\big[(t-t_{n})\big(\widehat{L}^{n}+{\cal E}^{n}\big)\big]\bar{\bm{e}}^{n}(t_{n})\Big]_{i},$
(SM50) $=\big[\big(\widehat{L}^{n}+\widehat{{\cal E}}^{n}\big)\bar{\bm{e}}^{n}\big]_{i},$
(SM51) $=-\frac{\alpha^{n}(x_{i+1/2})e^{n}(x_{i+1/2})-\alpha^{n}(x_{i-1/2})e^{n}(x_{i-1/2})}{h}.$
@@
We are now ready to present an error estimate for the fully discrete scheme (SM46). The steps in this proof track closely with those used in the proof of [15, Lemma 3], in which we developed a truncation error estimates for non-conservative, FD, Runge-Kutta discretizations of constant wave-speed linear advection problems.
Lemma SM6.3 (Truncation error estimate for $\Phi^{n}$).
@@
Assume that Footnote 10 holds, and let $\bar{\bm{e}}^{n}(t)$ be the exact solution of the local conservation law (SM42) at time $t\in[t_{n},t_{n+1}]$. Assume this function is sufficiently smooth. Then, the local truncation error of the one-step discretization (SM46) at time $t_{n+1}$ can be written as
(SM52) $\bm{\tau}^{n}$ $:=\bar{\bm{e}}^{n}(t_{n+1})-\Phi^{n}\bar{\bm{e}}^{n}(t_{n}),$
(SM53) $=\Big(-\delta t\widehat{{\cal E}}^{n}+\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n}\big)^{q+1}\Big)\bar{\bm{e}}^{n}(t_{n+1})+{\cal O}\Big(\big(\delta t\widehat{L}^{n}\big)^{q+2},\big(\delta t\widehat{{\cal E}}^{n}\big)^{2}\Big),$
@@
in which $\widehat{L}^{n}$ is the spatial discretization introduced in (SM43), and $\widehat{{\cal E}}^{n}$, as defined in Footnote 10, is the associated spatial discretization error.
Proof SM6.4.
Using (SM48), the solution to (SM42) at time $t=t_{n}$ may be written as
(SM54) $\bar{\bm{e}}^{n}(t_{n})=\exp\big(-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big)\bar{\bm{e}}^{n}(t_{n+1}).$
Plugging this into definition (SM52) gives
(SM55) $\bm{\tau}^{n}=\Big[I-\Phi^{n}\exp\big(-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big)\Big]\bar{\bm{e}}^{n}(t_{n+1}).$
@@
Next, plug in the series expansion for $\Phi^{n}$ given in (SM47) and simplify to give (abusing notation by dropping the vector $\bar{\bm{e}}^{n}(t_{n+1})$)
(SM56) $\begin{split}&I-\Phi^{n}\exp\big(-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big)\\ &=I-\Big[\exp\big(\delta t\widehat{L}^{n}\big)+\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n}\big)^{q+1}+{\cal O}\Big(\big(\delta t\widehat{L}^{n}\big)^{q+2}\Big)\Big]\exp\big(-\delta t(\widehat{L}^{n}+\widehat{{\cal E}}^{n})\big),\end{split}$
(SM57) $=\Big(\delta t\widehat{{\cal E}}^{n}-\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n}\big)^{q+1}\Big)+{\cal O}\Big(\big(\delta t\widehat{L}^{n}\big)^{q+2},\big(\delta t\widehat{{\cal E}}^{n}\big)^{2}\Big).$
Plugging this final expressing into (SM55) concludes the proof.
@@
From Lemma SM6.3 we want to develop a truncation error estimate for the ideal coarse-grid operator $\Psi_{\rm ideal}^{n,m\delta t}:=\prod_{j=0}^{m-1}\Phi^{n+j}$, i.e., $m$ fine-grid steps across the coarse-grid interval $t\in[t_{n},t_{n+m}]$. This can be developed rigorously using Taylor-series-type argument, but this is not necessary for our purposes since this analysis is heuristic anyway. Considering the form of (SM53) we suppose that the ideal coarse-grid operator satisfies the approximation
@@
(SM58) $\begin{split}&\bar{\bm{e}}^{n+m-1}(t_{n+m})-\Psi_{\rm ideal}^{n,m\delta t}\bar{\bm{e}}^{n}(t_{n})\approx\\ &\hskip 17.22217pt\Bigg[\sum_{j=0}^{m-1}\Big(\delta t\widehat{{\cal E}}^{n+j}-\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n+j}\big)^{q+1}\Big)\Bigg]\bar{\bm{e}}^{n+m-1}(t_{n+m}).\end{split}$
SM6.3 Error estimates for reconstructions
@@
We conclude Section SM6 with the following lemma: An error estimate for the left-shifted polynomial reconstructions that are described Sections 2.2 and SM1.
@@
Lemma SM6.5 (Error estimate for shifted linear reconstruction using $k$ cells).
@@
(SM59) $\big(\widetilde{R}^{\ell}\bar{\bm{u}}\big)_{i}$ $=u(x_{i-1/2})-\frac{(-h)^{k}\widetilde{\zeta}^{\ell}}{(k+1)!}\frac{\operatorname{d}\!^{k}u}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{i,\ell}(x_{i-1/2})},\quad\widetilde{\zeta}^{\ell}=(-1)^{\ell}\ell!(k-\ell)!,$
(SM60) $\big(R^{\ell}\bar{\bm{u}}\big)_{i}$ $=u(x_{i+1/2})-\frac{(-h)^{k}\zeta^{\ell}}{(k+1)!}\frac{\operatorname{d}\!^{k}u}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{i,\ell}(x_{i+1/2})},\quad\zeta^{\ell}=(-1)^{\ell+1}(\ell+1)!(k-\ell-1)!,$
@@
Proof SM6.6.
Because the reconstruction polynomial is the derivative of an interpolation polynomial (see Section SM1), this proof more-or-less follows the same steps as one used to derive the error for finite-difference approximations to a first derivative (these too are derivatives of interpolation polynomials); see e.g., [34, Section 8.1] or [15, Lemma 2].
Recall that for fixed $\ell$ that $\big(\widetilde{R}^{\ell}\bar{\bm{u}}\big)_{i}=q_{i}^{\ell}(x_{i-1/2})$ and $\big(R^{\ell}\bar{\bm{u}}\big)_{i}=q_{i}^{\ell}(x_{i+1/2})$, where $q_{i}^{\ell}(x)=\operatorname{d}\!Q_{i}^{\ell}/\operatorname{d}\!x$ is a degree at most k1k-1 polynomial recovering the cell averages of $e(x)$ over $k\geq 1$ cells $\{{\cal I}_{i-\ell},\ldots,{\cal I}_{i-\ell+k-1}\}$. Moreover, $Q_{i}^{\ell}$ is the unique polynomial of degree at most $k$ that interpolates $U(x):=\int^{x}u(\xi)\operatorname{d}\!\xi$ at the $k+1$ interfaces of the cells in the above stencil: $Q_{i}^{\ell}(x_{j+1/2})=U(x_{j+1/2})$ at points $\{x_{j+1/2}\}_{j=i-\ell-1}^{i-\ell-1-k}$.
@@
(SM61) $U(x)-Q^{\ell}(x)=\frac{K(x)}{(k+1)!}\frac{\operatorname{d}\!^{k+1}U}{\operatorname{d}\!x^{k+1}}\bigg|_{\xi_{\ell}(x)},\quad K(x):=\prod\limits_{j=-\ell-1}^{-\ell-1+k}(x-x_{j+1/2}),$
with $\xi_{\ell}(x)$ some (unknown) point in the interior of the interpolation stencil, $\xi_{\ell}(x)\in(x_{i-\ell-1/2},x_{i-\ell+k-1/2})$. Since $\operatorname{d}\!U/\operatorname{d}\!x=u$ and we desire an estimate for $u$, we differentiate (SM61) to get
(SM62) $u(x)-q^{\ell}(x)=\frac{1}{(k+1)!}\frac{\operatorname{d}\!K}{\operatorname{d}\!x}\bigg|_{x}\frac{\operatorname{d}\!^{k}u}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{\ell}(x)}+\frac{K(x)}{(k+1)!}\frac{\operatorname{d}\!^{k+1}u}{\operatorname{d}\!x^{k+1}}\bigg|_{\xi_{\ell}(x)}\frac{\operatorname{d}\!\xi_{\ell}}{\operatorname{d}\!x}\bigg|_{x}.$
To arrive at the estimates (SM59) and (SM60), all that is required is to evaluate (SM62) at $x=x_{-1/2}$ and $x=x_{1/2}$, respectively. To this end, notice from (SM61) that $K(x_{\pm 1/2})=0$, and, thus, the second added on the right-hand side of (SM62) vanishes when the estimate is evaluated at $x_{\pm 1/2}$. Next, from (SM61) we have that
(SM63) $\frac{\operatorname{d}\!K}{\operatorname{d}\!x}=\sum\limits_{s=-\ell-1}^{-\ell-1+k}\hskip 8.61108pt\prod\limits_{\begin{subarray}{c}j=-\ell-1\\ \neq s\end{subarray}}^{-\ell-1+k}(x-x_{j+1/2}).$
It is not difficult to show that $\operatorname{d}\!K(x_{-1/2})/\operatorname{d}\!x=(-h)^{k}\widetilde{\zeta}^{\ell}$, and $\operatorname{d}\!K(x_{1/2})/\operatorname{d}\!x=(-h)^{k}{\zeta}_{\ell}$. Evaluating (SM62) at $x_{\pm 1/2}$, and plugging in the above values concludes the proof.
SM7 MGRIT solution of linear conservation laws using standard, linear method-of-lines discretization